OpenStudy (anonymous):

What is the slope of the line between (3, –4) and (–2, 1)?

OpenStudy (anonymous):

i thought it was 2 but im not sure

OpenStudy (anonymous):

-2*

jimthompson5910 (jim_thompson5910):

Start by subtracting the y coordinates: 1 - (-4) = 1+4 = 5 Then subtract the x coordinates (in the same order): -2 - 3 = -5 Finally, divide the two differences (the y difference by the x difference): 5/(-5) = -1 So the slope is -1

OpenStudy (anonymous):

Whats the correct inequality for this graph?

OpenStudy (anonymous):

y\[\le\]3x-1

OpenStudy (anonymous):

are you sure?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Yes it is -1

OpenStudy (anonymous):

ok thanks!